OpenStudy (anonymous):

Solve by substitution. y = x − 1 2x − 3 y = negative-1 Infinite solutions (4, 3) (0, 0) (3, 4) @igreen

OpenStudy (igreen):

\(y = x − 1 \) \(2x − 3 y = -1\) Input x - 1 for y in the second equation: \(2x − 3 y = -1\) \(2x - 3(x - 1) = -1\) Distribute 3 into the parenthesis: \(2x - 3x - 3 = -1\) Subtract: \(-x - 3 = -1\) Add 3 to both sides: \(-x = 2\) Multiply -1 to both sides: \(x = -2\)

OpenStudy (igreen):

Now that we have the x value, we can plug it in any of the equations to find the y-value. \(y = x - 1\) Plug in -2: \(y = -2 - 1\) Subtract: \(y = -3\)

OpenStudy (anonymous):

oh someone beat us here

OpenStudy (anonymous):

UGH

OpenStudy (igreen):

Okay, I did something wrong..xD

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@R-Y-P